A general loses one sixth of his troops to a deadly virus. Of those that survive the virus, one fourth are killed in battle.
If there are 75 troops remaining, how many were present before the disease struck?

Answer by ikleyn(52777)

.
A general loses one sixth of his troops to a deadly virus. Of those that survive the virus, one fourth are killed in battle.
If there are 75 troops remaining, how many were present before the disease struck?
~~~~~~~~~~~~~~~~~~~~~~~~~

x - (x/6) - (1/4)*(x - (x/6)) = 75,   or

(5/6)x - (1/4)*(5/6)x = 75,  or

(1 - (1/4))*(5/6)x = 75,  or

(3/4)*(5/6)x = 75,  or

x = 75*(6/5)*(4/3) = 120.

Answer.  120 troops.